Circular No. 6598

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) K. Birkle and U. Thiele, Max Planck Institute for Astronomy, Hei-

delberg; and H. Böhnhardt, University Observatory, Munich, report: “A pattern of at least 12 straight dust striae appears almost identically on exposures with the Calar Alto Schmidt on Mar. 17.185 (see http://caserv. caha.es/CAHA/halebopp.html) and 18.172 UT at a nuclear distance of 1^{o}.5 to 3^{o} in the dust tail, the projected drift velocity of the pattern being ∼ 15 km/s. The striae are 1′ to 3′ wide, and some of them extend over > 1^{o} across the tail. The average position angle, ∼ 335^{o}, of the striae indicates that the pattern does not follow the synchronic direction for direct emission from the nucleus. The striae may thus be considered as due to secondary processes influencing the dust grains in the outer coma or in the dust-tail region. The overall appearance resembles the striae patterns seen in comets C/1975 V1 (West) and probably also in comet C/1910 A1 (Great January Comet). We also notice an enhanced brightness level in the dust-tail region northward of p.a. 315^{o}, this indicating enhanced dust activity in the nucleus after about Feb. 5.”

GRB 970111 AND GRB 970228 A. J. Castro-Tirado and J. Gorosabel, Laboratorio de Astrofísica Es-

pacial y Física Fundamental, Madrid, report, on behalf also of J. Heidt, T. Seitz, E. Thommes, C. Wolf, N. Lund, H. Pedersen, E. Costa, F. Frontera, J. Heise, J. in’t Zand, C. Bartolini, A. Guarnieri, A. Masetti, A. Piccioni and E. Palazzi: “Following the detection of GRB 970111 (IAUC 6533), deep frames (25 min in B and 10 min in R) were obtained on Jan. 12.2, Feb. 10.2 and 11.2 UT (only R on the last occasion) with the Calar Alto 2.2-m telescope (+ CAFOS). No fading object was detected, within the entire IPN/SAX error box (IAUC 6578), between the first set of images (taken 19 hr after the burst) and the later data. Any fading was < 0.2 mag for objects with B < 21, R < 20.8 and < 0.5 mag for those down to B = 23, R = 22.6. Taking into account the fact that GRB 970111 has a higher γ-ray fluence than GRB 970228 does, and unless the flux from x-ray to optical wavelengths was absorbed by intervening matter close to the source, the nondetection of either an x-ray counterpart (by the SAX narrow-field instruments) or a fading optical counterpart in the GRB 970111 field indicates a significant difference in the behavior of GRB 970111 in comparison to that of GRB 970228, for which a presumably-related x-ray/optical counterpart has been reported (IAUC 6576, 6584).”

1997 March 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden